I would appreciate any help on this problem, I've made up a table for this problem as the teacher had shown in class for other similar examples but my answer doesn't seem to check out.
"A Laboratory has 60 mL of a solution that is 40% HCl acid. How many mL of 15% HCL acid solution should be mixed with the 60 mL of 40% acid to obtain a solution of 25% HCl?"

Answer by stanbon(75887)

A Laboratory has 60 mL of a solution that is 40% HCl acid. How many mL of 15% HCL acid solution should be mixed with the 60 mL of 40% acid to obtain a solution of 25% HCl
-------------------
Active ingredient Equation:
0.40*60 + 0.15x = 0.25(x+60)
Multiply thru by 100 to get:
40*60 + 15x = 25x + 25*60
10 x = 15*60
x = 90 mL (amt. of 15% solution that must be added to the mixture)
=========================
Cheers,
Stan H.